OpenStudy (anonymous):

On the graph of the equation 8x – 3y = 24, what is the value of the y-intercept? Answer –8 8 –3 3

OpenStudy (aroub):

-3y=-8x+24 y=8/3x+8 y-intercept= 8

OpenStudy (amistre64):

the "y axis" IS the "x=0" axis the "y-intercept" IS the "x=0 intercept"

OpenStudy (anonymous):

so whats the answer because thats not it i have on top

OpenStudy (anonymous):

So in other words, if you make X=0, the value that you get for Y is what your intercept is.

OpenStudy (amistre64):

soo true..

OpenStudy (aroub):

or u can do it by return it to the form y=mx+b and b is the y-intercept

OpenStudy (anonymous):

David: Both amistre64 and aroub have shown you how to do it. One of them has even directly shown you what the answer is. Lol. Just try it.

OpenStudy (amistre64):

aroub has a "typo" tho :)

OpenStudy (anonymous):

There is indeed a typo.

OpenStudy (anonymous):

so its 8

OpenStudy (amistre64):

when x=0 we get: –3y = 24 ; what is the value for y?

OpenStudy (anonymous):

-8

OpenStudy (amistre64):

very good ;)

OpenStudy (anonymous):

so thats it?

OpenStudy (amistre64):

well, that all we NEED to do to it, anything else would just be a waste of time ....

OpenStudy (anonymous):

The reason it works is because the y intercepts are always in the form of (0,y), where y is ANY number. For example, if you have (0,100), that means the y intercept is 100. The 0 is always the value of the x. (x,y) ---> (0,y)